### v3

# Reconfigurable Four-Photon Interference in a Deployed Metropolitan Fiber Network

Abstract
Advanced quantum-networking protocols beyond point-to-point links rely critically on the ability to perform multi-photon interference under realistic operating conditions. Yet experimental validation of such higher-order interference effects in deployed metropolitan fiber networks remains limited. Here, we report a field demonstration of polarization-controlled reconfigurable four-photon interference over a commercially deployed metropolitan fiber network. Using a fully fiber-coupled linear-optical platform, we observe a fusion-type four-photon interference signature in the presence of real-world impairments, including photon loss, polarization drift, and timing uncertainty. By performing polarization-resolved measurements on two locally retained photons, we conditionally select distinct two-photon coincidence channels that exhibit Bell-like and N00N-like behavior under an idealized source model. Rather than pursuing multi-partite entanglement verification, this work focuses on establishing the technical feasibility of multi-photon interference and reconfigurable conditional state preparation in an operational network environment. These results serve as a systems-level validation milestone toward future multi-photon quantum-networking architectures that require robust interference performance outside the laboratory.
preprint: APS/123-QED
I Introduction
Quantum networks, with the ultimate vision of a quantum Internet, are poised to enable the transmission of quantum states and nonclassical correlations between spatially separated resources. Such capabilities are foundational for secure communication, distributed quantum computation, and networked quantum sensing beyond what is achievable with purely classical signals [1, 2]. Over the past 5 years, most quantum networking research on metro-scale fiber networks worldwide has emphasized the robust distribution and verification of bi-photon polarization entanglement over deployed telecommunication infrastructure, enabled by active polarization compensation, high-efficiency superconducting detection, and precise timing synchronization. For example, polarization entanglement from a type-II spontaneous parametric down-conversion (SPDC) source in the telecom O-band was distributed over 10 km of deployed fiber with active polarization stabilization in Singapore [3]. Researchers in Austria and Slovakia transmitted polarization-entangled SPDC photons across 248 km of deployed fiber, achieving sustained high visibility using automated polarization control and nonlocal dispersion compensation [4]. In New York City, polarization-entangled photon pairs at 1324 nm–-795 nm generated via four-wave mixing in warm rubidium vapor were distributed across a 34 km network with high fidelity and uptime [5]. In China, polarization-entangled photon pairs generated by a silicon photonic chip via spontaneous four-wave mixing were distributed over 30 km, with a CHSH Bell violation exceeding 27.8 $\sigma$ [6]. Additional metro-scale demonstrations have employed quantum dots for polarization-correlation distribution and related timing applications [7, 8]. Most significantly, researchers in Chicago employed polarization entanglement generated via type-0 cascaded second-harmonic generation SPDC to demonstrate quantum teleportation over 30.2 km of fiber co-propagating high-power classical telecom traffic [9], validating quantum–classical coexistence in shared infrastructure through optimized transmission bands and noise-mitigation techniques.
While these demonstrations establish that deployed fiber networks can support high-fidelity bi-photon entanglement distribution at metropolitan scales, extending such experiments to multi-photon interference—which underpins many advanced quantum-network protocols [10, 11, 12]—over deployed networks remains a significant challenge. Multi-photon experiments are substantially more sensitive to loss, background noise, polarization drift, and timing uncertainty, and are therefore often restricted to laboratory settings using fiber spools or short links [13, 14, 15, 16]. Nevertheless, developing and validating field-ready methods for multi-photon interference is an important intermediate step toward scalable quantum-network protocols that rely on higher-order correlations, multiplexing, and conditional heralding.
In this work, we demonstrate polarization-controlled reconfigurable multi-photon interference over a commercially deployed, metropolitan-scale telecom C-band fiber network. Using an all-fiber platform composed of a fully fiber-coupled bi-photon source and linear-optical fusion based on fiber beam splitters, we observe four-photon coincidence interference under temporal overlap. We use polarization-resolved local measurements as a reconfigurable control knob to select between distinct conditional coincidence channels with Bell-like and N00N-like behaviors associated with different multi-photon interference pathways. This functionality is presented as a network control and benchmarking primitive, enabling systematic exploration of photon loss sensitivity (with N00N states amplifying photon loss and distinguishability errors), polarization drift robustness, and interference stability in a deployed environment, rather than as a final application or a demonstration of entanglement distribution.
Beyond the specific experimental configuration, this work addresses a broader and largely untested assumption underlying many advanced quantum-network protocols [10, 11, 12] that multi-photon interference can be made to function reliably outside controlled laboratory environments. Although bi-photon entanglement distribution over deployed fiber has now been demonstrated by several groups mentioned above, substantially fewer experiments have examined the feasibility of higher-order interference processes in field conditions, despite their central role in protocols such as entanglement swapping, multiplexed repeaters, and multi-user-based protocols. Our results provide a direct field validation of multi-photon interference as a network-level resource, highlighting both its feasibility and its sensitivity to loss, noise, and source quality. A second key contribution of this work is the demonstration of reconfigurable conditional coincidence-channel selection in a deployed network. By performing polarization-resolved measurements on two locally retained photons, we switch between distinct four-photon coincidence channels corresponding to Bell-like and N00N-like correlation structure in the remaining two photons. This capability is presented as a network control primitive that enables systematic exploration of how different multi-photon interference channels behave under realistic fiber-network conditions.
The significance of this work is a field demonstration of a reconfigurable, polarization-controlled, multi-photon interference protocol on real-world telecommunication infrastructure in the presence of substantial loss and environmental perturbations. We explicitly focus on interference- and correlation-based signatures rather than entanglement verification, and we discuss the technical improvements required to enable rigorous multi-partite entanglement witnesses in future experiments.
The paper is arranged as follows. We begin with an overview of the deployed metropolitan fiber network infrastructure in downtown Chattanooga, followed by a theoretical model of the four-photon fusion stage and derive the polarization-dependent conditional coincidence structure corresponding to Bell-like and N00N-like correlation channels. We then describe the experimental methodology, including characterization of the fully fiber-coupled bi-photon source, synchronization of photon arrival times across the metro-scale network, and implementation of the fusion protocol. Next, we present field measurements demonstrating polarization-controlled reconfigurable four-photon interference and conditional remote two-photon coincidence signatures across spatially separated network nodes. Finally, we analyze the dominant performance limitations in the current field configuration and outline a pathway toward rigorous entanglement verification enabled by forthcoming upgrades to network hardware and source fidelity.
Figure 1: The EPB-IonQ Bohr-IV quantum network architecture.
Figure 2: A bird’s-eye view of the EPB-IonQ Bohr-IV quantum network architecture in downtown Chattanooga is shown in (a), with the segment utilized in this work highlighted in (b). Since Equipment Hubs A and B are located on Broad Street and Tenth Street, respectively, we label them as BQN and TQN. The deployed fiber loops between UTC-TQN and UTC-BQN are approximately 5 km and 10 km in length respectively.
II Overview of the deployed fiber network
EPB, a municipal electric utility and fiber-optic communication provider, and IonQ have partnered to operate a deployed quantum-network testbed in downtown Chattanooga. Designed to support photonic qubit generation, distribution, and measurement across multiple nodes, the network enables researchers from industry, government, and academia to test devices and protocols in a deployed fiber-optic environment. The network architecture, shown in Fig. 1(a), supports qubit communication over 4 dedicated fiber-optic strands and is therefore designated as the Bohr-IV Quantum Network. Further details of the architecture, including the Command Center, Equipment Hubs, and Quantum Nodes, are provided in Ref. [17]. Currently, seven metropolitan-scale quantum nodes have been deployed, including nodes owned by EPB, the University of Tennessee at Chattanooga (UTC), and IonQ, with planned expansion in the coming years as illustrated in Fig. 1. Each node is equipped with a quantum network interface console (QNIC), which provides physical connectivity and supports network reconfiguration. Through the QNIC, users can access real-time signal data via fiber-optic channels and interface with the network’s LAN layer to transmit configuration requests and retrieve diagnostics. Optimized for operation in the telecom C-band (1530–1570 nm), this reconfigurable infrastructure provides an ideal testbed for our field study of multi-photon interference and conditional projection over deployed fiber links. A bird’s-eye view of the EPB-IonQ Bohr-IV network is shown in Fig. 2(a), with the segment utilized in this work highlighted in Fig. 2(b).
III Theoretical Models
III.1 Four-Photon Fusion Stage and Conditional Projection Channels
In the type-II SPDC process used in this work, the nonlinear crystal produces photon pairs with orthogonal polarizations. For a single photon pair emitted into spatial modes $a$ and $b$, the two-photon component can be written in operator form as a superposition of orthogonally polarized emission terms,
$|\psi\rangle_{\text{pair}}\propto\left(a_{H}^{\dagger}b_{V}^{\dagger}+e^{i\phi}\,a_{V}^{\dagger}b_{H}^{\dagger}\right)|0\rangle,$ (1)
where the relative phase $\phi$ and the balance between the two terms depend on the source implementation and internal optics. In fully fiber-coupled, spatially degenerate sources, subsequent linear-optical processing and coincidence post-selection can select polarization-antisymmetric coincidence subspaces, which are exploited in the present work. In our fully fiber-coupled, spatially degenerate source, the two photons are separated probabilistically and routed into the interferometric network (Fig. 3). For the purposes of modeling the fusion stage, we represent each emitted pair as occupying two spatial modes with orthogonal polarizations and track how post-selected four-fold coincidences arise after interference at the final beam splitter.
As illustrated in Fig. 3, the idealized SPDC source (labeled ‘BPS’) generates two pairs of photons: one pair in modes $a,b$, and the other in modes $c,d$. These four photons are probabilistically split by the subsequent three fiber-coupled 50/50 beam splitters (labeled ‘BS’), with a success probability of 1/32. The resulting initial four-photon state, before arriving simultaneously at the final BS, is given by the tensor product:
$|\psi\rangle_{\text{initial}}=\frac{1}{2}\left(a_{H}^{\dagger}b_{V}^{\dagger}\pm a_{V}^{\dagger}b_{H}^{\dagger}\right)\left(c_{H}^{\dagger}d_{V}^{\dagger}\pm c_{V}^{\dagger}d_{H}^{\dagger}\right)|0\rangle$ (2)
$=\frac{1}{2}(|H_{a}V_{b}H_{c}V_{d}\rangle\pm|H_{a}V_{b}V_{c}H_{d}\rangle$
$\pm|V_{a}H_{b}H_{c}V_{d}\rangle+|V_{a}H_{b}V_{c}H_{d}\rangle),$
where the relative signs encode the polarization-correlation structure of the emitted pairs after the source’s internal optics and any post-selection used in the characterization stage. In the fusion stage, modes $a$ and $c$ are interfered on a 50/50 beam splitter to produce output modes $e$ and $f$, while modes $b$ and $d$ are retained for polarization-resolved conditioning.
A 50/50 BS transforms the creation operators of the input modes ($a,c$) into those of the output modes ($e,f$) according to:
$a_{p}^{\dagger}\to\frac{1}{\sqrt{2}}\left(e_{p}^{\dagger}+f_{p}^{\dagger}\right),\quad c_{p}^{\dagger}\to\frac{1}{\sqrt{2}}\left(e_{p}^{\dagger}-f_{p}^{\dagger}\right),$ (3)
where $p\equiv H,V$. Modes $b$ and $d$ are unaffected by the final BS.
Applying Eq. (3) to the four-photon component yields a superposition over output-number patterns in modes $e$ and $f$, including (i) coincidence terms with one photon in each output mode, and (ii) bunching terms with two photons in a single output mode. When the interfering photons are indistinguishable in all degrees of freedom, multi-photon interference suppresses particular coincidence pathways (Hong–Ou–Mandel–type suppression), producing a dip in appropriately chosen four-fold coincidence counts as the relative temporal delay is scanned. More explicitly, and assuming both input pairs are in the Bell singlet state, the resulting four-photon state can be written as follows:
$|\psi\rangle_{\text{final}}=\frac{1}{2\sqrt{2}}\Big[\left(|2_{H}\rangle_{e}-|2_{H}\rangle_{f}\right)|V_{b}V_{d}\rangle$ (4)
$-\frac{1}{\sqrt{2}}\left(|HV\rangle_{e}-|H\rangle_{e}|V\rangle_{f}+|V\rangle_{e}|H\rangle_{f}-|HV\rangle_{f}\right)|V_{b}H_{d}\rangle$
$-\frac{1}{\sqrt{2}}\left(|VH\rangle_{e}-|V\rangle_{e}|H\rangle_{f}+|H\rangle_{e}|V\rangle_{f}-|VH\rangle_{f}\right)|H_{b}V_{d}\rangle$
$+\left(|2_{V}\rangle_{e}-|2_{V}\rangle_{f}\right)|H_{b}H_{d}\rangle\Big].$
[…]
In this work, we use this polarization-controlled reconfigurable conditioning to switch between these two interference regimes in the deployed metropolitan fiber network and quantify the resulting changes in multi-photon coincidence statistics. Throughout, we refer to these outcomes as Bell-like and N00N-like conditional channels, and we reserve the term entanglement for future work that includes explicit multi-partite entanglement witnesses.
Figure 3: Schematic illustrating our scheme for interfering two photon pairs and implementing a four-photon fusion stage across three nodes, labeled UTC, TQN, and BQN, using a series of fiber-coupled BSs. Two photons are retained locally at UTC for polarization-resolved heralding, while two photons are transmitted through deployed fiber to the two remote nodes of BQN and TQN as shown in Fig. 2(b).
III.2 Polarization-Correlation Fringe Model for the Bell-like Channel
For completeness, we summarize the expected polarization-correlation dependence for the Bell-like conditional channel in the idealized case where a singlet-like two-photon component is selected. An antisymmetric two-photon polarization state with a relative phase delay $\Delta\phi$ between orthogonal polarizations can be written as
[…]
With
[…]
one obtains the idealized coincidence probabilities
$P_{HH,VV}=\frac{1}{4}(1-\cos\Delta\phi),$
$P_{HV,VH}=\frac{1}{4}(1+\cos\Delta\phi).$ (8)
In the present work, we use these expressions as a convenient reference model to compare against measured polarization-correlation fringes, while emphasizing that entanglement verification would require additional measurements (e.g., an entanglement witness or Bell test) beyond the scope of this field demonstration.
IV Experiments and Results
IV.1 Characterization of the Fully Fiber-Coupled Bi-Photon Source
Figure 4: Characterization of the fully fiber-coupled spatially-degenerate BPS through (a) HOM interference and (b) polarization-correlation fringe measurements with an LCVR, with the corresponding experimental results shown in (c) and (d), respectively. Photons are detected using SNSPDs, and two-photon coincidences are recorded with a coincidence window of 8 ns. The solid pink line in (c) is a Gaussian fit, while the solid green and pink lines in (d) are theoretical curves for the two-photon coincidence probabilities $P_{HH}$ and $P_{HV}$ respectively from Eq. (8) after mapping $\Delta\phi$ using the LCVR calibration data.
In this work, we employ an all-fiber-optic platform consisting of a continuous-wave (CW) fully fiber-coupled bi-photon source (manufactured by Qubitekk and labeled ‘BPS’) and fiber-coupled linear optics, including BSs, PBSs, and a variable delay line (VDL) (all manufactured by OZ Optics). The BPS is based on type-II SPDC in periodically-poled KTP crystal and is temperature-tuned to produce photon pairs near 1570 nm with a bandwidth of approximately 3 nm. Because the source output is fully fiber-coupled and spatially degenerate, the photons are separated and routed using fiber BSs as shown in Fig. 3.
We characterize the source using (i) Hong–Ou–Mandel (HOM) interference [18] and (ii) polarization-correlation fringe measurements based on the phase-dependent model in Eq. (8). The corresponding experimental schematics are shown in Figs. 4(a) and 4(b).
Throughout this work, the slow axis of the PM fiber connected to the reflected output of the PBS is rotated so that both output photons are $H$-polarized before detection by superconducting nanowire single-photon detectors (SNSPDs), maximizing detection efficiency. As shown in Fig. 4(a), temporal overlap at the BS is achieved by scanning the VDL. The characterization data in Figs. 4(c) and 4(d) were acquired using a two-photon coincidence window of 8 ns. The Gaussian fit in Fig. 4(c) yields a HOM visibility of 68 %, reported after subtracting accidental coincidences. The dominant sources of accidental coincidences arise from multi-pair emission in the CW SPDC process and background counts associated with imperfect polarization rejection and fiber-induced mode mixing. The curves in Fig. 4(d) are theoretical predictions from Eq. (8) after mapping $\Delta\phi$ using the LCVR wavelength–voltage calibration data provided in the Thorlabs datasheet (LCC1613-C) [19].
As shown in Fig. 4(d), the measured polarization-correlation fringe contrast is low, consistent with a substantial incoherent or mixed contribution in the source output. In our implementation, this manifests as an effective Bell-pair fraction of approximately $17~\%$ under the assumptions of the employed fringe model. We emphasize that the source is a commercial, closed, fully fiber-coupled unit with no available alignment degrees of freedom for performance optimization. The elevated background levels indicate significant contributions from undesired polarization components and imperfect mode indistinguishability. This source performance is the dominant limiting factor for the present field demonstration, primarily by reducing multi-photon interference contrast and four-fold coincidence rates.
We further attribute reductions in interference visibility to fiber-environment polarization sensitivity and imperfections in the all-fiber linear optics, including insertion loss at the VDL air gap, BS outputs imbalance, PBS’s PM fiber axis misalignment, and polarization mismatch across fiber mating sleeves. In the remainder of this paper, we therefore focus on robust, qualitative signatures of multi-photon interference and polarization-controlled reconfigurable conditional multi-state projection in the deployed network.
IV.2 Multi-Photon Coincidence Measurements Across the Network
[…]
The EPB–IonQ network introduces attenuation due to all-optical switching used for reconfigurable routing [17]: approximately 5 dB round-trip loss between UTC and TQN and approximately 10 dB round-trip loss between UTC and BQN. While such attenuation can be tolerated in some loss-robust networking protocols [20, 21], it substantially reduces four-fold coincidence rates in our multi-photon interference experiments. In the following sections we report the resulting acquisition times and discuss practical implications for scaling to higher-rate field demonstrations.
Figure 5: Characterization of the relative photon arrival times across the network. (a) Relative arrival times of the four photons, with one photon sent to TQN and another one to BQN, while the rest two are retained at UTC. (b)&(c) Zoomed-in views of the temporal correlation measurements between UTC–TQN and UTC–BQN, respectively. The relative delay between the two photons retained locally at UTC, represented by the purple line, is too small to be resolved in (a).
IV.3 Fusion Operation Across the Network
Figure 6: (a) Experimental layout for the four-photon fusion stage locally at UTC. (b) Experimental layout for the four-photon fusion stage across the network. The solid circles in (c)&(d) are the experimental results corresponding to (a)&(b), respectively. A dip in the four-fold coincidence counts indicates multi-photon interference at the fusion BS. (e) Four-fold polarization-resolved coincidence counts over the 16 measurement outcomes, recorded at the dip setting of (d). The solid pink lines in (c)&(d) are Gaussian fits.
With the source and arrival-time synchronization characterized, we demonstrate the four-photon fusion stage described in Section III.A across the network. To highlight network reconfigurability, in this experiment we retain the photons in modes $e$ and $f$ at UTC while routing modes $b$ and $d$ to BQN and TQN, respectively (Fig. 6(b)). We first perform the fusion operation locally by retaining all four photons at UTC (Fig. 6(a)). The local and network fusion results are shown in Figs. 6(c) and 6(d), respectively. As temporal overlap is achieved by adjusting the VDL, multi-photon interference suppresses distinguishable-path coincidence contributions, producing a dip in the measured four-fold coincidence rate. We interpret these dips as evidence of successful temporal overlap and indistinguishability sufficient to observe fusion-stage interference in both local and deployed-fiber configurations.
To further characterize the polarization-dependent structure of the four-fold coincidence events at the network-fusion setting, we fix the VDL at the dip position in Fig. 6(d) and record four-fold coincidences across all $2^{4}=16$ polarization outcomes (Fig. 6(e)). The corresponding four-photon coincidence counts are shown in Fig. 6(e), with four pronounced peaks for $H_{b}H_{f}V_{e}V_{d}$, $V_{b}V_{f}H_{e}H_{d}$, $H_{b}V_{f}H_{e}V_{d}$, and $V_{b}H_{f}V_{e}H_{d}$, confirming the presence of the four-photon ($e,f,b,d$) components in Eq. (4). These measurements confirm the presence of the expected four-photon coincidence components in the post-selected detection subspace and serve as a consistency check for subsequent polarization-controlled reconfigurable conditional measurements.
IV.4 Bell-like Conditional Channel Across the Network
Figure 7: (a) Experimental layout for polarization-controlled generation of the Bell-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. Data shown in (c)&(d) correspond to (a)&(b), respectively. The solid green and pink circles in (c)&(d) represent the measured four-photon coincidence counts $C_{H_{e}H_{f}}\big|_{V_{b}H_{d}}$ and $C_{H_{e}V_{f}}\big|_{V_{b}H_{d}}$, respectively. The thick solid green and pink curves in panels (c)&(d) denote the theoretical curves for the coincidence probabilities $P_{HH}$ and $P_{HV}$ from Eq. (8), respectively, while the thinner green and pink lines are provided as visual guides for the experimental data.
We next demonstrate polarization-controlled reconfigurable conditional projection corresponding to a Bell-like coincidence channel. We retain the two conditioning photons in modes $b$ and $d$ locally at UTC and route modes $e$ and $f$ through deployed fiber loops to TQN and BQN, respectively, before returning for detection. By projecting the locally retained photons $b$ and $d$ into cross-polarized outcomes (e.g., $V_{b}H_{d}$ or $H_{b}V_{d}$) using the PBSs highlighted in cyan in Figs. 7(a) and 7(b), and by recording four-fold coincidences, we select a conditional subset of remote events whose polarization-correlation dependence follows the theoretical model in Eq. (8). The local and deployed-network configurations are shown in Figs. 7(a) and 7(b), with the corresponding results presented in Figs. 7(c) and 7(d).
For the data shown, the conditioning two photons in modes $b$ and $d$ are detected first, defining the conditional subset for the subsequent detections associated with the two photons traversing the deployed fiber loops. Although the network data are noisier due to environmental disturbances and loss, the overall polarization-dependent four-fold coincidence trends remain visible. We emphasize that these measurements do not exceed standard inseparability thresholds and are therefore presented as polarization-controlled Bell-like conditional correlation signatures on a deployed network.
IV.5 N00N-like Conditional Channel Across the Network
Figure 8: (a) Experimental layout for polarization-controlled operation of the N00N-like conditional channel locally at UTC. (b) Experimental layout for the operation across the network. Data shown in (c)&(d) corresponds to (a)&(b), respectively. The solid green circles in (c)&(d) represent the measured four-photon coincidence counts with photons $b$ and $d$ projected into the same-polarization modes of $H_{b}H_{d}$. The thick solid pink lines in (c)&(d) are Gaussian fits, while the thinner green lines are provided as visual guides for the experimental data.
[…]
The observed visibility is suboptimal, with elevated dip values, consistent with the source and interferometric limitations discussed in Section IV.A. Also note that the most direct way to quantitatively characterize a N00N state is by measuring its Heisenberg-limited phase sensitivity [22, 23]. While our current network configuration and count rates do not allow such a measurement within reasonable acquisition times, this will be the focus of a dedicated future study.
V Conclusion and Discussion
In summary, we have demonstrated polarization-controlled reconfigurable four-photon interference and conditional two-photon projection over a metropolitan-scale deployed fiber network using an all-fiber linear-optical platform. By synchronizing photon arrival times across the deployed network and performing polarization-resolved conditioning on locally retained photons, we observe distinct four-fold coincidence interference signatures corresponding to Bell-like and N00N-like conditional channels. These results extend multi-photon interference experiments beyond the laboratory and into a real-world telecommunication infrastructure characterized by substantial loss, polarization drift, and environmental perturbations.
From a systems perspective, this work serves two primary roles. First, it provides a field validation of multi-photon interference as a viable network-level resource. Many advanced quantum-network protocols explicitly rely on the ability to generate and interfere multiple photons with sufficient indistinguishability [10, 11, 12], yet experimental evidence for such processes in deployed fiber has remained limited. Our results show that four-photon interference signatures can be observed on a metropolitan network, while also quantifying the practical limitations imposed by source infidelity, fiber loss, and timing uncertainty.
Second, the polarization-controlled switching between Bell-like and N00N-like conditional coincidence channels demonstrates a programmable network control primitive based entirely on linear optics and local measurements. This capability enables systematic benchmarking of deployed fiber links under different correlation structures and provides a pathway toward adaptive quantum network operation, where conditional multi-photon processes can be selected dynamically in response to network conditions. Notably, these capabilities remain meaningful even in the absence of entanglement verification, making them directly relevant to near-term quantum network development.
The current performance limitations are dominated by (i) source quality (limited polarization-correlation contrast and reduced indistinguishability) and (ii) network loss introduced by reconfigurable all-optical switching [17]. Both limitations directly impact multi-photon interference visibility and four-fold coincidence rates. Near-term improvements include replacing the present spatially degenerate commercial BPS with a higher-fidelity source (commercial or custom Sagnac-based [24]) and reducing photon-routing loss through planned all-optical switching upgrades. With improved source fidelity and network transmission, future work will implement explicit entanglement witnesses for the conditional channels demonstrated here.
In this work, all photons are ultimately detected at UTC after traversing the deployed fiber loops. Extending to true multi-node operation, with coincidence detection at physically separated nodes and independently running clocks, will require tighter timing synchronization (e.g., via the White Rabbit protocol) [25, 26], and a network configuration supporting distributed time-tag correlation without centralized detection. Finally, integrating automatic polarization control protocols (e.g., as in Ref. [27]) will be essential for long-duration stability of multi-photon interference in deployed fiber environments. These developments form the foundation for our next stage of experiments on the EPB-IonQ Bohr-IV quantum network.
acknowledgments
JEH, AFE, GSA, and TL acknowledge support from the U.S. National Science Foundation (NSF) through the ExpandQISE program under Award No. 2426699. JEH, AFE, and TL also acknowledge support from the NSF CCSS program under Award No. 2503630. KR and TL acknowledge support from the U.S. National Institute of Standards and Technology (NIST) through the CIPP program under Award No. 60NANB24D218. MMH and GS acknowledge support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, through the Quantum Internet to Accelerate Scientific Discovery Program under Field Work Proposal No. 3ERKJ381, and from the NSF under Award No. DGE-2152168.